OpenStudy (anonymous):

An investor invested a total of $1300 in two mutual funds. One fund earned a 7% profit value while the other earned 4% profit. If the investor's profit was $85, how much was in each mutual fund? The amount invested in the fund that earned 7% was __________. The amount invested in the fund that earned 4% was ________>

OpenStudy (anonymous):

please help i dont understand what to do

OpenStudy (anonymous):

525 and 775 i got for answers

OpenStudy (wolf1728):

Those are not the answers. I'm working on equations to try and help you.

OpenStudy (anonymous):

ok ty

OpenStudy (wolf1728):

First equation is pretty simple x - amount invested in 7% account y - amount invested in 4% account x + y = 1300

OpenStudy (anonymous):

a+b=1300 .07a+.04b=85

OpenStudy (wolf1728):

Gee Amy that is impressive!! Here's what I came up with x = Amount invested in 7% interest account y = Amount invested in 4% interest account A) x + y = $1300 B) .07x + .04y = $85 Multiply equation B by -25 yields -1.75x -1y = -2,125 Adding this equation to equation B yields -.75x = -825 x = 1,100

OpenStudy (anonymous):

1,100 is 7% profit and 825 is 4% profit

OpenStudy (anonymous):

that does not add up

OpenStudy (wolf1728):

No the 825 is only part of that equation. I did not post the y value. if x + y = 1300 and x = 1100 what does y equal?

OpenStudy (anonymous):

unless 200 goes for 4% and 1100 goes for 7% i hate math and my next class is the second part to this math and have not done this is 13 years

OpenStudy (wolf1728):

1100 * .07 = $77 200 * .04 = $8 77 + 8 = $85 :-)

OpenStudy (anonymous):

so 77 is 7% 8 is 4 %

OpenStudy (wolf1728):

Yes

OpenStudy (wolf1728):

There's one thing you should know and that's how to solve for 2 equations You did very well in coming up with one equation but I think you need help in solving those.

OpenStudy (anonymous):

yeah its been a while it took me over three days and 6 hours to do one over and over again

OpenStudy (wolf1728):

Here is part of y solution: A) x + y = $1300 B) .07x + .04y = $85 Multiply equation B by -25 yields -1.75x -1y = -2,125 Do you know why I multiplied equation B by -25?

OpenStudy (anonymous):

you have to multiply both sides to get the equation

OpenStudy (anonymous):

Use elimination method

OpenStudy (wolf1728):

Yes, you always have to do the same thing to both sides of the equation. Yes, I was using the elimination method. You do remember huh?

OpenStudy (anonymous):

it just getting it in the right order

OpenStudy (wolf1728):

Yes, I was messing up a lot on this one. Adding -1.75x to 1 - that sort of thing.

OpenStudy (anonymous):

yeah i have another one

OpenStudy (wolf1728):

okay

OpenStudy (anonymous):

an investor invested a total of $2800 in two mutual funds. One fund earned a 6% profit value while the other earned 2% profit. If the investor's profit was $120, how much was in each mutual fund? The amount invested in the fund that earned 6% was----. The amount invested in the fund that earned--- 2% was

OpenStudy (anonymous):

yay more i hate these

OpenStudy (anonymous):

but how do you know how to eliminate it with what

OpenStudy (wolf1728):

A) x = y = $2,800 B) .06x + .02y = $120

OpenStudy (anonymous):

x+y=2800 .06+.02y=120

OpenStudy (anonymous):

now elimnation part i get wrong all the time

OpenStudy (anonymous):

would it be 12 or 8

OpenStudy (wolf1728):

Okay, we have to eliminate an x or a y from the equations when they are added together. (What do you mean by 12 or 8?)

OpenStudy (anonymous):

i added them sorry

OpenStudy (wolf1728):

Before you add them you have to determine which variable is to be eliminated. So let's suppose BEFORE we add equation A & B - what variable do we eliminate?

OpenStudy (anonymous):

x i think

OpenStudy (wolf1728):

Either can be eliminated but I think it is easier to eliminate y. To eliminate 'y' we multiply equation B times -50 (If we wanted to elimnate 'x' we'd have to multiply by equation B by -16.6666666667 - not very neat is it?)

OpenStudy (anonymous):

nope

OpenStudy (wolf1728):

And if you are wondering where the -50 comes from, it is the amount needed to produce a minus y in Equation B

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

300 and 100 i got

OpenStudy (wolf1728):

.06x + .02y = $120 So, multiplying Equation B by -50 produces -3x -y = -6,000

OpenStudy (anonymous):

im confused now

OpenStudy (anonymous):

the other one the answer was 1100 and 200

OpenStudy (wolf1728):

Yess 1100 and 200 were the previous answers.

OpenStudy (wolf1728):

Okay - this a place where you should NOT get confused. We just multiplied equation B by -50 to get -3x -y = -6,000 Now what do we do next?

OpenStudy (anonymous):

Divide

OpenStudy (wolf1728):

No - think about it. In order to solve for 2 unknowns we need 2 equations.

OpenStudy (anonymous):

A) x + y = $2800 B) .06x + .02y = $120 Multiply equation B by -50 yields -3x -y = -6,000

OpenStudy (wolf1728):

That's right so now that we have that "new" equation B what do we do with it?

OpenStudy (anonymous):

Use elimination method

OpenStudy (wolf1728):

Yes, and how do we eliminate a variable?

OpenStudy (wolf1728):

We multiplied Equation B times -50. Do you remember why?

OpenStudy (anonymous):

to eliminate b

OpenStudy (wolf1728):

or to eliminate y (to use my nomenclature)

OpenStudy (anonymous):

ok so now we eliminate x

OpenStudy (wolf1728):

well we have to eliminate y - you rememer what to do?

OpenStudy (anonymous):

Adding this equation to equation B yields -.50x = 6000

OpenStudy (wolf1728):

Actually we have to add the modified equation B to equation A x = y = $2,800 -3x -y = -6,000

OpenStudy (anonymous):

-2x-y=3200 is what i got unless i am doing it wrong

OpenStudy (wolf1728):

Yes, you shouldn't have any y value (and the 3200 should be negative). So, it should be -2x = -3200

OpenStudy (anonymous):

-2x=-3200

OpenStudy (wolf1728):

and so x = 1600

OpenStudy (anonymous):

1200 is y

OpenStudy (wolf1728):

yes that's right

OpenStudy (anonymous):

6% is 1600 and 2% would be 1200

OpenStudy (wolf1728):

yes

OpenStudy (anonymous):

thank you I am not gonna do this assignment again i got the passing grade thank you for all your help and time

OpenStudy (wolf1728):

okay - you know what would help? Learning to solve for 2 unknowns with an easy example.

OpenStudy (wolf1728):

For example, try solving this: 3X + 2Y = 20 X + Y = 9